OpenStudy (anonymous):

Determine the equation of the line tangent to g(x)=2x^4+5/x-3 at x=-1 Write your answer in slope-intercept form (y = mx + b).

OpenStudy (anonymous):

f'(x)= 8x^3+5x^-1

OpenStudy (anonymous):

Sorry dont know how to explain

OpenStudy (anonymous):

take the derivative replace x by -1 to get your slope replace x by -1 in origninal function to get the second coordinate of the point use the point slope formula to find the equation

OpenStudy (anonymous):

f'(x)=8x^3-5

OpenStudy (anonymous):

not clear if this is \[g(x)=2x^4+\frac{5}{x-3}\] or \[g(x)=2x^4+\frac{5}{x}-3\] but i am guessing it is the first one

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

ok then \[g(-1)=2-5-3=-6\] so your point is (-1,-6)

OpenStudy (anonymous):

and \[g'(x)=8x^3-\frac{5}{x^2}\] therefore \[g'(-1)=-8-5=-13\] so you have the point and the slope, should be good from there yes?

OpenStudy (anonymous):

wait how did you get 8x^3-5/x^2?

OpenStudy (anonymous):

and how do you solve for b?